### v1

@@ Abstract
In this work, we investigate the discovery reach of a new physics model, the Inert Doublet Model, at an $e^{+}e^{-}$ machine with centre-of-mass energies $\sqrt{s}$ of 240 and 365 GeV. Within this model, four additional scalar bosons ($H,\,A,\,H^{+}$ and $H^{-}$) are predicted. Due to an additional symmetry, the lightest new scalar, here chosen to be $H$, is stable and provides an adequate dark matter candidate. The search for pair production of the new scalars is investigated in final states with two electrons or two muons, in the context of the future circular collider proposal, FCC-ee. Building on previous studies in the context of the CLIC proposal, this analysis extends the search to detector-level objects, using a parametric neural network to enhance the signal contributions over the Standard Model backgrounds, and sets limits in the $M_{A}-M_{H}$ vs $M_{H}$ plane. With a total integrated luminosity of 10.8 (2.7)  ab-1 for $\sqrt{s}=$240 (365) GeV, almost the entire phase-space available in the $M_{A}-M_{H}$ vs $M_{H}$ plane is expected to be excluded at 95% CL, reaching up to $M_{H}$$=110\,(165)$GeV. The discovery reach is also explored, reaching $M_{H}$$=108\,(157)\,\text{G\text{eV}}$ for $M_{A}-M_{H}$$=15$GeV  at $\sqrt{s}=$240 (365) GeV.
@@ 1 Introduction
Despite the numerous cosmological observations pointing towards the existence of dark matter (DM) ParticleDataGroup:2024cfk, the specific nature of DM is still an ever elusive mystery. This lack of understanding is further compounded by the fact that the abundance of DM is predicted to significantly outweigh that of normal matter. Assuming DM can interact directly with normal matter, it should be possible to detect it here on Earth. This can happen in a number of ways: DM passing through our solar system could directly interact with particles, imparting energy that can be detected in direct detection experiments; DM particles could modify Standard Model (SM) interactions and be detected indirectly through precision measurements; or, DM can be produced at a collider, leading for example to a missing energy signature. More details can be found in Ref. ParticleDataGroup:2024cfk.
We consider the case in which DM particles are weakly interacting massive particles. A particle of this nature, however, is not contained in the SM, and as such alternatives theories — or extensions to the SM — are required to include this. In the context of extensions to the SM such as two-Higgs-doublet models, it is possible, via the addition of an exact $\mathbb{Z}_{2}$ symmetry, to make the lightest scalar boson from the additional doublet stable, and hence a suitable DM candidate. The study presented here considers a search for one such theory: the Inert Doublet Model (IDM) Deshpande:1977rw,Barbieri:2006dq,Cao:2007rm,IDMrunIRecast,IDMbenchmarks.
Although many phenomenological studies exist, dedicated studies of the IDM have not yet been performed by the experimental collaborations. Existing analyses are generally not sensitive to a large part of the parameter space due to too high requirements on the missing transverse energy, and their recast leads to very loose constraints (see e.g. discussion in Dercks:2018wch). In this paper, we present a search for the pair production of the extra scalars in $e^{+}e^{-}$ collisions at centre-of-mass energies ($\sqrt{s}$) of 240 and 365 GeV  in the context of the future circular collider of CERN, FCC-ee FCC:2018evy,Blondel:2021ema,Agapov:2022bhm, using a final state with a pair of same-flavour leptons (electrons or muons).
@@ 2 The Inert Doublet Model
$\mathbb{Z}_{2}:\quad\phi_{D}\longrightarrow-\phi_{D},\quad\quad\phi_{S}\longrightarrow\phi_{S},\quad\quad SM\longrightarrow SM.$ (1)
@@
$\begin{split}V=-\frac{1}{2}[m^{2}_{11}(\phi_{S}^{\dagger}\phi_{S})+m^{2}_{22}(\phi_{D}^{\dagger}\phi_{D})]+\frac{\lambda_{1}}{2}(\phi_{S}^{\dagger}\phi_{S})^{2}+\frac{\lambda_{2}}{2}(\phi_{D}^{\dagger}\phi_{D})^{2}\\ \quad+\lambda_{3}(\phi_{S}^{\dagger}\phi_{S})(\phi_{D}^{\dagger}\phi_{D})+\lambda_{4}(\phi_{S}^{\dagger}\phi_{D})(\phi_{D}^{\dagger}\phi_{S})+\frac{\lambda_{5}}{2}[(\phi_{S}^{\dagger}\phi_{D})^{2}+(\phi_{D}^{\dagger}\phi_{S})^{2}].\end{split}$ (2)
@@
where $v$ and $M_{h}$ are the vacuum expectation value (vev) and mass stemming from the SM-like doublet and are fixed through experimental measurements to $v\,\sim\,246\,\text{G\text{eV}}$ and $M_{h}\,\sim\,125\,\text{G\text{eV}}$, respectively. $H,\,A,$ and $H^{\pm}$ denote the novel so-called dark scalars, and $\lambda_{i}$ are couplings in the potential with $\lambda_{345}\equiv\lambda_{3}+\lambda_{4}+\lambda_{5}$. Here, $\lambda_{2}$ is the coupling constant for the quartic vertex between four IDM particles, and $\lambda_{345}$ is the coupling constant for the vertex between IDM particles and the SM Higgs boson. The couplings between the new scalars and the SM gauge bosons are determined solely by quantities from the SM electroweak sector. It should be noted that the scalars from the two doublets do not mix due to the $\mathbb{Z}_{2}$ symmetry, therefore there are no additional mixing parameters.
@@
The $\mathbb{Z}_{2}$ symmetry leads to multiple interesting phenomenological features. First, the symmetry forbids any direct coupling between the IDM particles and the SM fermion fields, and as such the IDM particles only couple directly to bosons. Second, the symmetry requires that all terms in the Lagrangian involving IDM fields have an even number of IDM fields (2 or 4). As the second doublet does not acquire a vacuum expectation value, the symmetry remains exact and IDM particles are always produced in pairs. Finally, it also causes the lightest of the IDM particles to be stable and therefore a suitable DM candidate [^1] A detailed discussion on regions where the IDM can render the observed relic density can be found in e.g. Kalinowski:2020rmb.. In principle, any of the additional scalars can serve as a DM candidate. We here chose $H$, which automatically implies a mass hierarchy $M_{A}$,$M_{H^{\pm}}$$>$$M_{H}$.
@@
Given that the IDM particles do not couple to the SM fermions, a typical signature at a collider is that of SM electroweak gauge boson plus missing energy, where the latter stems from the DM candidates. In addition, the model can also render monojet and monophoton signatures Belyaev:2018ext,Belyaev:2016lok as well as processes containing the invisible decay of the 125 GeV SM-like scalar. This analysis focusses on the same-flavour dilepton (electrons or muons) plus missing energy final state, with example production mechanisms shown in figure 1.
@@
Figure 1: Leading-order Feynman diagrams of: (left) $HH\ell^{+}\ell^{-}$ production, (middle) $HH\ell^{+}\ell^{-}\nu\bar{\nu}$ production, and (right) $\mathup{{{Z}}}h\rightarrow HH\ell^{+}\ell^{-}$ production, with $\ell=e,\mu,\tau$.
@@
The IDM is subject to a large number of theoretical and experimental constraints. In this work, we follow the scan presented in Ref. Ilnicka:2015jba, with more recent updates presented in Refs Ilnicka:2018def,Kalinowski:2018ylg,Dercks:2018wch,Kalinowski:2020rmb,Braathen:2024lyl [^2] Note we do not apply the two-loop constraints on the scalar couplings presented in Braathen:2024lyl, but here chose to use the leading-order values for constraints.. In particular, we include constraints on the potential stemming from vacuum stability, perturbativity and perturbative unitarity, as well as electroweak precision observables (EWPO), evaluated through the publicly available tool 2HDMC Eriksson:2009ws,Eriksson:2010zzb. Results for the electroweak precision constraints are tested using the oblique parameters $S,\,T,\,U$ Altarelli:1990zd, Peskin:1990zt,Peskin:1991sw,Maksymyk:1993zm and comparing to the latest PDG values ParticleDataGroup:2024cfk. We also require that the parameter point resides in the inert vacuum Ginzburg:2010wa. We furthermore test for the most recent bound on SM Higgs boson decays to invisible particles in cases where this decay is allowed onshell, $\text{BR}\left(h\,\rightarrow\,\text{inv}\right)\,\leq\,0.11$ ATLAS:2023tkt. Regarding dark matter constraints, we calculate the predictions using Micromegas Belanger:2018ccd and require agreement with the upper bound for relic density as measured by the Planck experiment Planck:2018vyg. Direct detection constraints are taken from the LuxZeplin collaboration LZ:2024zvo. Finally, we also make sure the mass hierarchy obeys the constraints from a recast of searches for supersymmetric neutralinos at LEP reinterpreted within the IDM Lundstrom:2008ai (referred to as LEP SUSY recast in the following), and that the particles would not contribute to the decay widths of the electroweak gauge bosons at leading order. In the remainder of this document, we will label points as allowed that fulfill all above constraints, if not mentioned otherwise. More details on the scan setup and the applied constraints can be found in the references above.
@@
A set of benchmark points (BPs) were evaluated in Ref. Kalinowski:2018ylg to identify representative regions of the phase space where the production is still allowed by existing constraints. For the analysis targeting only the same-flavour dilepton final state, the signal cross section is usually dominated by the $AH$ production, and the sensitivity will depend mostly on $M_{H}$ and $M_{A}-M_{H}$. Instead of just estimating the sensitivity on the benchmark points, we adopt a strategy to scan the parameter space in $M_{H}$ and $M_{A}-M_{H}$, with several scenarios described below to study the dependency on the other parameters: $M_{H^{\pm}}$, $\lambda_{2}$ and $\lambda_{345}$.
@@
Diagrams involving quartic couplings between IDM scalar bosons do not participate to the production of the final state under study. Hence, there is no sensitivity to the $\lambda_{2}$ coupling, which is set to 0.1 everywhere. Concerning diagrams involving the coupling $\lambda_{345}$ of the new scalars with the SM $h$ (see Fig. 1 right) two scenarios are considered: one where $\lambda_{345}=0$ and one where $\lambda_{345}$ is set to a relatively large value estimated using a linear parametrisation in $M_{H}$, $\lambda_{345}=0.0018\,\times\,\left(M_{H}-72\,\text{G\text{eV}}\right)$. This parametrisation is chosen to approach the maximal value allowed given recent results on direct detection searches for dark matter. As DM bounds in addition depend on the mass of the unstable neutral scalar, $M_{A}$, the actual maximal value might be lower than the one given by this parametrisation.
@@
Similarly, we investigate two scenarios that differ in the choice of $M_{H^{\pm}}$: one with $M_{H^{\pm}}$$=$$M_{A}$, and one with $M_{H^{\pm}}$--$M_{A}$set to be within the maximum value allowed. Maximal mass splittings are in general mainly constrained from requiring perturbative unitarity as well as electroweak precision data [^3] We thank J. Braathen for useful discussions regarding this point.. Note, however, that the latter also depends on the absolute mass scale. In order to estimate the maximally allowed mass differences, we first perform a random scan where we fix $\lambda_{2}$ but let all other parameters float. The resulting allowed parameter space is shown in figure 2. The points that have been simulated according to the fixed mass differences discussed above are displayed in green. As the absolute scale also plays a role in the determination of the oblique parameters, some points we chose, although overlaying with the allowed regions of a general scan, are forbidden by EWPO for specific choices of mass differences. These are displayed in red, but we chose to keep them for the final result for consistency reasons.
[…]
To summarise, three different scenarios are considered, where we vary $M_{H}$ and $M_{A}$, and set $\lambda_{345}$ and $M_{H^{\pm}}$to specific values:
  • (S1) $M_{H^{\pm}}$$=$$M_{A}$ and $\lambda_{345}=0$;
  • (S2) $M_{H^{\pm}}$$=$$M_{A}$ and $\lambda_{345}$ close to the maximum value allowed, referred to as $\lambda_{max}$; this scenario is meant to test the impact of production through the SM $h$ (see Fig. 1 right);
  • (S3) within maximum $M_{H^{\pm}}$, referred to as $M_{H^{\pm}}^{max}$, and $\lambda_{max}$; this scenario is meant to test the contribution of the $H^{+}H^{-}$ production (see Fig. 1 middle).

### v2

@@ Abstract
We present a search for new scalar bosons predicted by the Inert Doublet Model at an $e^{+}e^{-}$ machine with centre-of-mass energies $\sqrt{s}$ of 240 and 365 GeV. Within this model, four additional scalar bosons ($H,\,A,\,H^{+}$ and $H^{-}$) are predicted. Due to an additional symmetry, the lightest new scalar, here chosen to be $H$, is stable and provides an adequate dark matter candidate. The search for pair production of the new scalars is investigated in final states with two electrons or two muons, in the context of the future circular collider proposal, FCC-ee. Building on previous studies in the context of the CLIC proposal, this analysis extends the search to detector-level objects, using a parametric neural network to enhance the signal contributions over the Standard Model backgrounds, and sets projected exclusion and discovery contours in the $M_{A}-M_{H}$ vs. $M_{H}$ plane. With a total integrated luminosity of 10.8 (2.7)  ab-1 for $\sqrt{s}=$240 (365) GeV, the discovery reach for the model goes up to $M_{H}$$108\,(157)\,\text{G\text{eV}}$ for $M_{A}-M_{H}$=15=15GeV at $\sqrt{s}=$240 (365) GeV. For exclusion, almost the entire phase-space available in the $M_{A}-M_{H}$ vs $M_{H}$ plane is expected to be ruled out at 95% CL, reaching up to $M_{H}$=110(165)=110\,(165)GeV.
@@ 1 Introduction
Despite the numerous cosmological observations pointing towards the existence of dark matter (DM) [1], the specific nature of DM is still an ever elusive mystery. This lack of understanding is further compounded by the fact that the abundance of DM is predicted to significantly outweigh that of normal matter. Assuming DM can interact directly with normal matter, it should be possible to detect it here on Earth. This can happen in a number of ways: DM passing through our solar system could directly interact with particles, imparting energy that can be detected in direct detection experiments; DM particles could modify Standard Model (SM) interactions and be detected indirectly through precision measurements; or, DM can be produced at a collider, leading for example to a missing energy signature. More details can be found in Ref. [1].
We consider the case in which DM particles are weakly interacting massive particles. A particle of this nature, however, is not contained in the SM, and as such alternatives theories — or extensions to the SM — are required to include this. In the context of extensions to the SM such as two-Higgs-doublet models, it is possible, via the addition of an exact $\mathbb{Z}_{2}$ symmetry, to make the lightest scalar boson from the additional doublet stable, and hence a suitable DM candidate. The study presented here considers a search for one such theory: the Inert Doublet Model (IDM) [2, 3, 4, 5, 6].
Although many phenomenological studies exist, dedicated studies of the IDM have not yet been performed by the experimental collaborations. Existing analyses are generally not sensitive to a large part of the parameter space due to too high requirements on the missing transverse energy, and their recast leads to very loose constraints (see e.g. discussion in [7]). In this paper, we present a search for the pair production of the extra scalars in $e^{+}e^{-}$ collisions at centre-of-mass energies ($\sqrt{s}$) of 240 and 365 GeV  in the context of the future circular collider of CERN, FCC-ee [8, 9, 10, 11]. Within the IDM, the main target process is the pair-production $e^{+}e^{-}\,\rightarrow\,AH$, where $A$ and $H$ are additional new physics scalars and $H$ serves as the dark matter candidate. The scalar $A$ subsequently decays into ZHZ\,H, and we use a final state with one same-flavour, opposite-sign lepton pair (electrons or muons) accompanied by missing energy from the invisible dark scalars.
Similar studies for this channel have already been performed in previous work within the CLIC environment, where higher centre-of-mass energies are also investigated, see e.g. [12, 13]. In this work, we extend on the previous studies by using full reconstruction. We also extend the interpretation to include the full kinematically-available mass ranges in a systematic way, thanks to the use of a parametric neural network.
@@ 2 The Inert Doublet Model
$\mathbb{Z}_{2}:\,\phi_{D}\longrightarrow-\phi_{D},\;\phi_{S}\longrightarrow\phi_{S},\;SM\longrightarrow SM.$ (1)
@@
$V$ $=$ $-\frac{1}{2}[m^{2}_{11}(\phi_{S}^{\dagger}\phi_{S})+m^{2}_{22}(\phi_{D}^{\dagger}\phi_{D})]+\frac{\lambda_{1}}{2}(\phi_{S}^{\dagger}\phi_{S})^{2}$
$+\frac{\lambda_{2}}{2}(\phi_{D}^{\dagger}\phi_{D})^{2}+\lambda_{3}(\phi_{S}^{\dagger}\phi_{S})(\phi_{D}^{\dagger}\phi_{D})$
$+\lambda_{4}(\phi_{S}^{\dagger}\phi_{D})(\phi_{D}^{\dagger}\phi_{S})+\frac{\lambda_{5}}{2}[(\phi_{S}^{\dagger}\phi_{D})^{2}+(\phi_{D}^{\dagger}\phi_{S})^{2}].$
@@
where $v$ and $M_{h}$ are the vacuum expectation value (vev) and mass stemming from the SM-like doublet and are fixed through experimental measurements to $v\,\sim\,246\,\text{G\text{eV}}$ and $M_{h}\,\sim\,125\,\text{G\text{eV}}$, respectively. H,A,H,\,A, and $H^{\pm}$ denote the novel so-called dark scalars, and $\lambda_{i}$ are couplings in the potential with $\lambda_{345}\equiv\lambda_{3}+\lambda_{4}+\lambda_{5}$. Here, $\lambda_{2}$ is the coupling constant for the quartic vertex between four IDM particles, and $\lambda_{345}$ is the coupling constant for the vertex between IDM particles and the SM Higgs boson. The couplings between the new scalars and the SM gauge bosons are determined solely by quantities from the SM electroweak sector. It should be noted that the scalars from the two doublets do not mix due to the $\mathbb{Z}_{2}$ symmetry, therefore there are no additional mixing parameters.
@@
The $\mathbb{Z}_{2}$ symmetry leads to multiple interesting phenomenological features. First, the symmetry forbids any direct coupling between the IDM particles and the SM fermion fields, and as such the IDM particles only couple directly to bosons. Second, the symmetry requires that all terms in the Lagrangian involving IDM fields have an even number of IDM fields (2 or 4). As the second doublet does not acquire a vacuum expectation value, the symmetry remains exact and IDM particles are always produced in pairs. Finally, it also causes the lightest of the IDM particles to be stable and therefore a suitable DM candidate [^1] A detailed discussion on regions where the IDM can render the observed relic density can be found in e.g. [5].. In principle, any of the additional scalars can serve as a DM candidate. We here chose $H$, which automatically implies a mass hierarchy $M_{A}$$M_{H^{\pm}}$$M_{H}$.
@@
Given that the IDM particles do not couple to the SM fermions, a typical signature at a collider is that of SM electroweak gauge boson plus missing energy, where the latter stems from the DM candidates. In addition, the model can also render monojet and monophoton signatures [14, 15] as well as processes containing the invisible decay of the 125 GeV SM-like scalar. This analysis focusses on the same-flavour dilepton (electrons or muons) plus missing energy final state, with example production mechanisms shown in figure 1.
@@
Figure 1: Leading-order Feynman diagrams of: (top) $HH\ell^{+}\ell^{-}$ production, (middle) $HH\ell^{+}\ell^{-}\nu\bar{\nu}$ production, and (bottom) $Zh\rightarrow HH\ell^{+}\ell^{-}$ production, with $\ell=e,\mu,\tau$.
@@
The IDM is subject to a large number of theoretical and experimental constraints. In this work, we follow the scan presented in Ref. [6], with more recent updates presented in Refs [16, 17, 7, 5, 18] [^2] Note we do not apply the two-loop constraints on the scalar couplings presented in [18], but here chose to use the leading-order values for constraints.. In particular, we include constraints on the potential stemming from vacuum stability, perturbativity and perturbative unitarity, as well as electroweak precision observables (EWPO), evaluated through the publicly available tool 2HDMC [19, 20]. Results for the electroweak precision constraints are tested using the oblique parameters S,T,US,\,T,\,U [21, 22, 23, 24] and comparing to the latest PDG values [1]. We also require that the parameter point resides in the inert vacuum [25]. We furthermore test for the most recent bound on SM Higgs boson decays to invisible particles in cases where this decay is allowed onshell, $\text{BR}\left(h\,\rightarrow\,\text{inv}\right)\,\leq\,0.11$ [26]. Regarding dark matter constraints, we calculate the predictions using Micromegas [27] and require agreement with the upper bound for relic density as measured by the Planck experiment [28]. Direct detection constraints are taken from the LUX-ZEPLIN collaboration [29]. Finally, we also make sure the mass hierarchy obeys the constraints from a recast of searches for supersymmetric neutralinos at LEP reinterpreted within the IDM [30] (referred to as LEP SUSY recast in the following), and that the particles would not contribute to the decay widths of the electroweak gauge bosons at leading order. In the remainder of this document, we will label points as allowed that fulfill all above constraints, if not mentioned otherwise. More details on the scan setup and the applied constraints can be found in the references above.
@@
A set of benchmark points (BPs) were evaluated in Ref. [17] to identify representative regions of the phase space where the production is still allowed by existing constraints. For the analysis targeting only the same-flavour dilepton final state, the signal cross section is usually dominated by the $AH$ production, and the sensitivity will depend mostly on $M_{H}$ and $M_{A}-M_{H}$. Instead of just estimating the sensitivity on the benchmark points, we adopt a strategy to scan the parameter space in $M_{H}$ and $M_{A}-M_{H}$, with several scenarios described below, and summarised in Table 1, to study the dependency on the other parameters: $M_{H^{\pm}}$, $\lambda_{2}$ and $\lambda_{345}$.
Scenario $M_{H^{\pm}}$ $\lambda_{345}$ motivation
S1 $M_{A}$ 0 targets main contribution process
S2 $M_{A}$ $\lambda_{max}$ switches on $h$-Strahlung process
S3 $M_{H^{\pm}}^{max}$ $\lambda_{max}$ switches on contributions from $H^{+}\,H^{-}$ production
Table 1:
Scenarios considered in this work for the scan in the $\left(M_{H},M_{A}-M_{H}\right)$ plane. Parameters are chosen such that they subsequently add the contributions of the diagrams in figure 1 to the signal process. See text for details.
@@
Diagrams involving quartic couplings between IDM scalar bosons do not participate to the production of the final state under study at leading order. Hence, there is no sensitivity to the $\lambda_{2}$ coupling, which is set to 0.1 everywhere. Concerning diagrams involving the coupling $\lambda_{345}$ of the new scalars with the SM $h$ (see Fig. 1 bottom) two scenarios are considered: one where $\lambda_{345}=0$ and one where $\lambda_{345}$ is set to a relatively large value, referred to as $\lambda_{max}$, estimated using a linear parametrisation in $M_{H}$, $\lambda_{max}=0.0018\,\times\,\left(M_{H}-72\,\text{G\text{eV}}\right)$. This parametrization is chosen as an envelope of the parameter points in the $\left(M_{H},\,\lambda_{345}\right)$ plane that were allowed by early LUX-ZEPLIN results for direct detection [31] and approaches the maximal value for the coupling combination allowed by these constraints. As DM bounds in addition depend on the mass of the unstable neutral scalar, $M_{A}$, the actual maximal value might be lower than the one given by this parametrisation.
@@
Next, we investigate a scenario that differs in the choice of $M_{H^{\pm}}$, from $M_{H^{\pm}}$$M_{A}$ to $M_{H^{\pm}}$-$M_{A}$  set to be within the maximum value allowed, giving a value of $M_{H^{\pm}}$ referred to as $M_{H^{\pm}}^{max}$. Maximal mass splittings are in general mainly constrained from requiring perturbative unitarity as well as electroweak precision data [^3] We thank J. Braathen for useful discussions regarding this point.. Note, however, that the latter also depends on the absolute mass scale. In order to estimate the maximally allowed mass differences, we first perform a random scan where we fix $\lambda_{2}$ but let all other parameters float. The resulting allowed parameter space is shown in figure 2. The points that have been simulated according to the fixed mass differences discussed above are displayed in green. As the absolute scale also plays a role in the determination of the oblique parameters, some points we chose, although overlaying with the allowed regions of a general scan, are forbidden by EWPO for specific choices of mass differences. These are displayed in red, but we chose to keep them for the final result for consistency reasons. Note that $M_{H^{\pm}}-M_{A}\neq\,0$ is only realized in one of the three scenarios we consider; therefore, the points that would be forbidden here are not explicitly labeled in our final result plots.
Figure 2: Allowed points in a general scan with $\lambda_{2}\,=\,0.1$ (purple, ++), as well as points used in the simulation (green, x). Some of the latter are forbidden by electroweak precision observable constraints, which were implemented via the oblique parameters S,T,US,\,T,\,U (red, *).